OpenStudy (anonymous):

Solving Systems by Elimination. 2x - 3y = 61 2x + y = -7

OpenStudy (anonymous):

So, 2x - 3y = 61 - 2x + y = -7 -------------- 0x + (-4y) = 68 -4y = 68 y = 17

OpenStudy (anonymous):

Agh. Sorry, I'm just stressed and trying to calm.. I had a hard day today..y = 17 *Hides face*

OpenStudy (anonymous):

-17*

OpenStudy (anonymous):

Okay, night..

OpenStudy (anonymous):

I'm just gonna do this tomorrow with my teacher, and do my Biology now.